# M3-Net: A Cost-Effective Graph-Free MLP-Based Model for Traffic Prediction

### v2

@@ Abstract.
Achieving accurate traffic prediction is a fundamental but crucial task in the development of current intelligent transportation systems. The main challenges of the existing deep learning approaches are that they either depend on a complete traffic network structure or require intricate model designs to capture complex spatio-temporal dependencies. These limitations pose significant challenges for the efficient deployment and operation of deep learning models on large-scale datasets. To address these challenges, we propose a cost-effective graph-free Multilayer Perceptron (MLP) based model M3-Net for traffic prediction. Extensive experiments conducted on multiple real datasets demonstrate the superiority of the proposed model in terms of prediction performance and lightweight deployment.
Traffic prediction, spatio-temporal modeling, multilayer perceptron, mixture of experts
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11doi: 10.1145/XXXXXX.XXXXXXccs: Information systems Spatial-temporal systemsisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. INTRODUCTION
Recently, although there has been significant growth of works for urban traffic prediction as well as analogous ST forecasting tasks (Jin et al., 2023a; Liu et al., 2023b), the aforementioned challenges are still not fully addressed. Some methods focus on modeling ST correlations by utilizing a single shared model across all locations (Sherstinsky, 2020; Yu et al., 2019; Lai et al., 2018). Such approaches often overlook the inherent heterogeneity among different regions and fail to explicitly capture diverse spatial-temporal interaction patterns, making it difficult for the model to adapt to varying local characteristics. Another line of research builds graph-structured representations to enhance spatial modeling (Yu et al., 2018; Li et al., 2018; Wu et al., 2019a; Jin et al., 2022; Guo et al., 2019; Jin et al., 2023b; Li et al., 2022), typically propagating information along static or learned graphs. While effective in representing spatial relations, these models heavily rely on predefined topologies and incur significant computational overhead, which limits scalability in dynamic or large-scale networks. More recently, attention-based methods have been developed to learn long-range ST dependencies in a fully data-driven manner without explicit graph construction (Vaswani et al., 2017; Xu et al., 2020; Liu et al., 2023a). Although these methods improve modeling flexibility, they still struggle to distinguish between local heterogeneity and global homogeneity and often suffer from high memory and computational costs, hindering their deployment in real-world traffic systems.
@@ 3.1. Experimental Setup
We evaluate our model on four normal datasets and three large-scale datasets. The normal datasets include PEMS03, PEMS04, PEMS07 and PEMS08, Which are collected from California highways (Song et al., 2020). In the experiments, we use the traffic flow of the last 12 time steps to predict the traffic flow of the next 12 time steps, and record the prediction performance of the 3rd, 6th, 12th steps and the average. Our model is implemented by Pytorch 1.7 with vGPU-32GB. We set the number of spatial groups as 10 in spatial MLP, the number of experts in channel MLP as 4 and the number of M3-Layers as 3 by default. The dimension of hidden representations in our model is set as 128 and the dimension of embedding layer is set as 32. We set the Adam optimizer with an initial learning rate of 0.002, where the learning rate follows a step-wise decay strategy, and the batch size is set as 64.
@@ 3.2. Performance Evaluation
We compare our proposed model with fourteen state-of-art baselines as follows: DCRNN (Li et al., 2018), STGCN (Yu et al., 2018), GWNet (Wu et al., 2019b), AGCRN (Bai et al., 2020), StemGNN (Cao et al., 2020), GMAN (Zheng et al., 2020), MTGNN (Wu et al., 2020), DGCRN (Li et al., 2023), STNorm (Deng et al., 2021) and STID (Shao et al., 2022). The evaluation metrics are mean absolute errors (MAE), root mean squared errors (RMSE) and mean absolute percentage errors (MAPE) averaged over five times for one hour ahead prediction. The comparison results of our proposed model with other baselines on normal datasets are shown in Table 2. We can observe that our model achieves the best results on most metrics across multiple datasets and also achieves the second-best results on the remaining metrics. Baselines such as GWNet, MTGNN, AGCRN, StemGNN and DGCRN are commonly used state-of-the-art models based on adaptive graphs in the field of ST prediction. The overall performance of M3-Net surpasses these models, demonstrating that the combination of the spatio-temporal embeddings and a finely designed MLP architecture can outperform some more complexly designed models.

### v3

@@ Abstract.
Achieving accurate traffic prediction is a fundamental but crucial task in the development of current intelligent transportation systems. The main challenges of the existing deep learning approaches are that they either depend on a complete traffic network structure or require intricate model designs to capture complex spatio-temporal dependencies. These limitations pose significant challenges for the efficient deployment and operation of deep learning models on large-scale datasets. To address these challenges, we propose a cost-effective graph-free Multilayer Perceptron (MLP) based model M3-Net for traffic prediction. Extensive experiments conducted on multiple real datasets demonstrate the superiority of the proposed model in terms of prediction performance and lightweight deployment. Our code is available at https://github.com/jinguangyin/M3_NET
Traffic prediction, spatio-temporal modeling, multilayer perceptron
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11doi: 10.1145/3746252.3760815ccs: Information systems Spatial-temporal systemsisbn: 979-8-4007-2040-6/2025/11
@@ 1. INTRODUCTION
Recently, although there has been significant growth of works for urban traffic prediction as well as analogous ST forecasting tasks (jin2023spatio; liu2023largest), the aforementioned challenges are still not fully addressed. Some methods focus on modeling ST correlations by utilizing a single shared model across all locations (RNN; lstm; lai2018modeling). Such approaches often overlook the inherent heterogeneity among different regions and fail to explicitly capture diverse spatial-temporal interaction patterns, making it difficult for the model to adapt to varying local characteristics. Another line of research builds graph-structured representations to enhance spatial modeling (stgcn; dcrnn; graphwavenet; jin2022automated; astgcn; jin2023stnpp; li2022automated), typically propagating information along static or learned graphs. While effective in representing spatial relations, these models heavily rely on predefined topologies and incur significant computational overhead, which limits scalability in dynamic or large-scale networks. More recently, attention-based methods have been developed to learn long-range ST dependencies in a fully data-driven manner without explicit graph construction (attention; sttn; staeformer). Although these methods improve modeling flexibility, they still struggle to distinguish between local heterogeneity and global homogeneity and often suffer from high memory and computational costs, hindering their deployment in real-world traffic systems.
@@ 3.1. Experimental Setup
We evaluate our model on four normal datasets and three large-scale datasets. The normal datasets include PEMS03, PEMS04, PEMS07 and PEMS08, Which are collected from California highways (song2020spatial). In the experiments, we use the traffic flow of the last 12 time steps to predict the traffic flow of the next 12 time steps, and record the prediction performance of the 3rd, 6th, 12th steps and the average. Our model is implemented by Pytorch 1.7 with vGPU-32GB. We set the number of spatial groups as 10 in spatial MLP, the number of experts in channel MLP as 4 and the number of M3-Layers as 3 by default. The dimension of hidden representations in our model is set as 128 and the dimension of embedding layer is set as 32. We set the Adam optimizer with an initial learning rate of 0.002, where the learning rate follows a step-wise decay strategy, and the batch size is set as 64.
@@ 3.2. Performance Evaluation
We compare our proposed model with fourteen state-of-art baselines as follows: DCRNN (dcrnn), STGCN (stgcn), GWNet (gwn), AGCRN (agcrn), StemGNN (stemgnn), GMAN (gman), MTGNN (mtgnn), DGCRN (dgcrn), STNorm (stnorm) and STID (stid). The evaluation metrics are mean absolute errors (MAE), root mean squared errors (RMSE) and mean absolute percentage errors (MAPE) averaged over five times for one hour ahead prediction. The comparison results of our proposed model with other baselines on normal datasets are shown in Table 2. We can observe that our model achieves the best results on most metrics across multiple datasets and also achieves the second-best results on the remaining metrics. Baselines such as GWNet, MTGNN, AGCRN, StemGNN and DGCRN are commonly used state-of-the-art models based on adaptive graphs in the field of ST prediction. The overall performance of M3-Net surpasses these models, demonstrating that the combination of the spatio-temporal embeddings and a finely designed MLP architecture can outperform some more complexly designed models.